Raymond Mining Corporation has 9.1 million shares of common stock outstanding, 350,000 shares of 4% $100 par value preferred stock outstanding, and 155,000 7.50% semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $39 per share and has a beta of 1.55, the preferred stock currently sells for $95 per share, and the bonds have 10 years to maturity and sell for 110% of par. The market risk premium is 7.9%, T-bills are yielding 5%, and Raymond Mining’s tax is 40%.

a. What is the firm’s market value capital structure?

Market value
Debt ?
Equity _. ?
Preferred stock

  ?

b. If Raymond Mining is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 3 decimal places.)

Discount rate ______ %
Answer #1

a.Debt Market Value =Par Value *110%*Number of Bonds =1000*110%*155000=170,500,000
b. Equity Value =Number of Shares*Share Price =9,100,000*39 =354,900,000
c. Preferred Stock =Number of Preferred Stocks*Share Price =350000*95 =33,250,000

b.
Par Value =1000
Semi annual coupon =7.5%*1000/2 =37.50
Number of periods =10*2 =20
Price =110%*1000 =1100
YTM using excel formula =2*RATE(20,37.50,-1100,1000) =6.1466%
Cost of Preferred Stock =Dividend/Price =4%*100/95 =4.21053%
Cost of Equity =Risk free rate+Beta*(Market Return-Risk free rate) =5%+1.55*7.9% =17.245%
Total Value =170,500,000+354,900,000+33,250,000 =558650000
Cost of Capital =Weight of Equity*Cost of equity+Weight of Preferred Stock*Cost of Preferred Stock+Weight of Debt*Cost of Debt*(1-Tax Rate)
=354,900,000/558650000*17.245%+33,250,000/558650000*4.21053%+170,500,000/558650000*6.466%*(1-40%) =12.332%